# Machine Learning Should Maximize Welfare, Not (Only) Accuracy

### v1

@@ 2.1 Welfare economics: crash course
Welfare economics makes explicit the idea that individuals have agency. Intuitively, this states that individuals (i) want things, (ii) know things, and (iii) act – to get what they want, using what they know. These notions are formally accounted for by modeling utility functions (want), private information (know), and decision-making, e.g., rational or behavioral (act). Any policy that aims to advance welfare must take these into account. Often this requires the planner to make additional efforts, such as to elicit preferences, create incentives for truthful reporting, or infer how users will respond to different policy choices. These are challenging, but give the planner power: if incentives can be aligned, then it becomes possible to harness the willingness of users to invest effort for improving outcomes for all; consider public goods and services, crowdfunding platforms, open-source software, and collaborative knowledge bases.
2.2 Connections to machine learning.
We believe that machine learning has much to gain from adopting a welfare perspective: when inputs represent humans, it becomes possible to promote overall social benefit (efficiency), and imperative to consider its distribution across the population (equity). Current tools already push forward on these ideas, but only to a limited extent. One reason is that standard learning objectives are notoriously underspecified; this has implications on e.g. robustness (D’Amour et al., 2022), explainability (see Rudin, 2019), and fairness (e.g., Rodolfa et al., 2020; Coston et al., 2021; Black et al., 2022). But underspecification also presents an opportunity for steering outcomes toward socially beneficial states. Consider how the idea of model multiplicity (Breiman, 2001; Marx et al., 2020; Hsu and Calmon, 2022), i.e., that there is typically a large set of (approximately) optimal models, connects to the notion of an efficient Pareto front. The challenge lies in how to provide a social planner effective means to choose a preferred operating point. In some cases, this will be diffucult; in others, it may be achievable with tools as simple as regularization—if chosen appropriately (Levanon and Rosenfeld, 2021). In terms of limited resources, machine learning already offers many relevant tools, such as constraints on cardinality (e.g., top-$k$ prediction) or error rates (e.g., precision and recall). Also relevant are tools from cost-sensitive (e.g., Elkan, 2001) and decision-focused learning (Mandi et al., 2024). But to be effective, these must be adapted to account for agency: how users report information, how they act in response to the learned classifier, and how they interact with each other. This idea will be key to our framework, as discussed next.
3 Three Orders of Welfare-Maximizing ML
@@
Our framework presents a hierarchy of problem formulations for welfare-maximizing learning, organized into three ‘orders’ of gradually increasing complexity. The orders are built in a bottom-up fashion: The lowest order (Order 0) coincides with the standard objective of maximizing predictive accuracy, but provides a novel welfare perspective suited for social tasks. Each higher order then builds on and generalizes the one below it, this by adding to the objective another layer of economic complexity: first focusing on system decisions (Order 1), and then on user choices (Order 2). As such, our framework is based on accuracy maximization at its core, but enables—and in fact requires—to explicitly model resources, allocations, and agency. It also requires to specify the role predictions play in shaping social outcomes. An illustration of the framework and its orders is given in Fig. 1.
We begin with a welfare interpretation of supervised learning, and then proceed to discuss the three framework orders. Further prospects and challenges are discussed in Appx. 5.
Supervised learning from a welfare perspective.
Given labeled training data of pairs $(x,y)$ sampled iid from an unknown distribution ${D}$, the goal in supervised learning is to find a function $h$ from a chosen class $H$ whose predictions $\hat{y}=h(x)$ are accurate on future examples in expectation:
Since in our setting examples $(x,y)$ are associated with individuals in the population ${D}$, then from a welfare perspective, we can think of Eq. (2) as prescribing a particular social welfare function—a special case of Eq. (1)—in which: (i) utility to users derives from accurate predictions, (ii) all users share the same utility function, $u(x,y;{\pi}_{h})=\mathds{1}{\{{y=h(x)}\}}$, (iii) possible outcomes include correct and incorrect predictions, hence (iv) the policy is degenerate, $u(x,y;{\pi}_{h})=u(x,y;h)$), and finally (v) weights are uniform, $w(x)\equiv 1$.
From accuracy to welfare maximization.
Despite an apparent ‘neutrality’, the above reveals that accuracy maximization in fact makes an (implicit) statement about social preferences. This entails a particular notion of equity—one that derives operationally from the predictive task at hand, rather than from a planned or designated social goal. But learning is of course not bound to maximizing objectives only of the form of Eq. (2). If our true goal is to maximize welfare, then we should modify the objective to support this goal. For example, if users differ in how much they benefit from accurate predictions, then we can plug in an appropriate utility function $u$; or if there are exogenous constraints on predictions, then we can express these via the policy ${\pi}_{h}$. Such steps are certainly possible, but introduce challenges beyond those found in standard machine learning tasks. Towards this, we pave a path that gradually transforms Eq. (2) to support increasingly complex economic considerations.
@@ 3.1 Order 0: Allocating Accuracy
Accuracy will be beneficial to users in any a platform or service that relies on predictions; consider recommendation systems, online market platforms, financial aid tools, or diagnistic health services. We argue that in such cases, and as long as learning is prone to some level of error (which is plausible), then accuracy itself is a limited resource—simply because not all users can obtain the same level of accuracy. Thus, the choice of classifier becomes, in effect, a decision on which users will be ‘allocated’ more accurate predictions, and which less. This then begs the question: who should be allocated an accurate prediction? One answer is users who most need it, or who benefit mostly. In the reasonable case where there is natural variation in individual utilities, we can to optimize a utility-weighted accuracy objective:
@@
Maximizing welfare through learning presents new algorithmic challenges. One task is to identify learners that optimize a given social welfare function—intrinsically, a question of identifying a desirable point on the Pareto front. Another task is to generate all points on the frontier (see, e.g., Navon et al., 2021). One possible approach is to cast Eq. (3) as a problem of distribution shift. Here shift can stem from several factors, including the welfare function weights, user utilities, and user inputs. But agency means that utilities can be misreported and features manipulated; as a result, the way in which the distribution shifts becomes dependent on the choice of classifier, indirectly through how users respond. This makes Eq. (3) a challenging instance of model-dependent distribution shift (Drusvyatskiy and Xiao, 2023). Here the interaction between the system and its users can be modeled as a Stackelberg game (Chen et al., 2020). This perspective has been useful for questions on learnability, e.g. via generalization bounds that rely on strategic VC analysis (Zhang and Conitzer, 2021; Sundaram et al., 2023) or that accommodate feature-dependent utilities (Pardeshi et al., 2024). Nonetheless, many important open questions remain.
@@ 3.2 Order 1: Incorporating System Decisions
Generally, predictions are useful for the system if they aid in making better decisions about uncertain outcomes. Since decisions in social settings often directly prescribe how to allocate limited resources to users, a useful next step is to encode them explicitly into the objective. We formalize this idea by modeling a system that makes decisions about users (e.g., who to hire) through predictions (e.g., resume screening). Consider a user $x$ with label $y$, and denote by $a$ the action the system takes (e.g., hire or not). Let ${r}(a,y)$ be the reward for the system on this user given action $a$ (e.g., the quality of the candidate, if hired). The direct dependence of ${r}$ on $y$ means that if we know $y$, then we can write the optimal action as $a^{*}={\pi}(y)$ for some policy ${\pi}$. Since $y$ is generally unknown, the common approach is to replace $y$ in ${\pi}$ with a prediction ${\hat{y}}=h(x)$, denoted $a={\pi}(h(x))={\pi}_{h}(x)$, in hopes that better predictions translate to better decisions. Such policies, referred to as prediction policies (Kleinberg et al., 2015), are appropriate when the uncertainty in $y$ is a stronger factor for outcomes than potential causal effects (we discuss this distinction further in Appendix 5.4). In this setting, choosing to work with prediction policies gives reason for the system to optimize $h$ for accuracy (Eq. (2)).

### v2

@@ 2 Welfare economics: crash course
Welfare economics makes explicit the idea that individuals have agency. Intuitively, this states that individuals (i) want things, (ii) know things, and (iii) act – to get what they want, using what they know. These notions are formally accounted for by modeling utility functions (want), private information (know), and decision-making, either rational or behavioral (act). Any policy that aims to advance welfare must take these into account. Often this requires the planner to make additional efforts, such as to elicit preferences, create incentives for truthful reporting, or infer how users will respond to different policy choices. These are challenging, but give the planner power: if incentives can be aligned, then it becomes possible to harness the willingness of users to invest effort individually for improving outcomes for all; consider public goods and services, crowdfunding platforms, open-source software, and collaborative knowledge bases.
3 Towards welfare-maximizing machine learning
We believe that machine learning has much to gain from adopting a welfare perspective: when inputs represent humans, it becomes possible to promote overall social benefit (efficiency), and imperative to consider how benefits are distributed across the population (equity). We next exemplify this idea on the most basic task of supervised classification, interpreted from a welfare perspective.
Accuracy as social welfare.
Consider a supervised learning task with data sampled as $(x,y)\sim{D}$, and in which inputs $x$ describe individuals (e.g., loan application, resume, movie viewing history, past consumer choices), labels $y$ as describing outcomes of interest (e.g., loan return or default, job qualification, movie rating, satisfaction), and ${D}$ represents the population. The goal is to find a function $h$ from a chosen class $H$ whose predictions $\hat{y}=h(x)$ are accurate on future examples in expectation:
Because pairs $(x,y)$ are associated with individuals in a population, accuracy admits a welfare interpretation. In particular, it prescribes a specific social welfare function—a special case of Eq. (1)—in which: (i) utility to users derives from accurate predictions, (ii) all users share the same utility function, $u(x,y;{\pi}_{h})=\mathds{1}{\{{y=h(x)}\}}$, (iii) possible outcomes include correct and incorrect predictions, hence (iv) the policy is degenerate, $u(x,y;{\pi}_{h})=u(x,y;h)$, and finally (v) weights are uniform, $w(x)\equiv 1$.
Despite an apparent ‘neutrality’, the above reveals that accuracy maximization in fact makes a particular implicit statement about social preferences. This entails a specific notion of equity—one that derives operationally from the predictive task at hand, rather than from a planned or designated social agenda. But learning is of course not bound to maximizing objectives only of the form of expected accuracy. If our true goal is to maximize welfare, then we should modify the objective to support this goal. For example, if users differ in how much they benefit from accurate predictions, then we can plug in an appropriate utility function $u$; or if there are exogenous constraints on predictions, then we can express these via the policy ${\pi}_{h}$. We will dive deeper into these possibilities in Sec. 4.
3.1 Why welfare?
The above exercise conveys a clear message: once we learn, we are inevitably making some statement about welfare outcomes. This idea applies broadly, including to today’s generative models. A primary benefit of augmenting learning frameworks with a well-defined welfare notion is that such statements can be made explicit. This would not only be an option, but a requirement—working within a welfare-aware framework necessitates making precise statements about efficiency and equity, and makes clear who it is making these statements (e.g., the learner, platform, or government). It also requires to articulate the core components: namely resources, allocation, agency, and policy. Aside from these conceptual benefits, an integrated welfare formalism offers several practical benefits:
Structure.
Welfare economics relies on well-defined social welfare functions, and casts the problem of promoting well-being as one of optimizing such functions. Welfare functions that depend on predictions can naturally be interpreted as learning objectives; in fact, some already are, as we argue in Sec. 4. Given that machine learning has come to excel at optimizing complex objectives over expressive function classes, welfare objectives provide a new target for applying these practices.
Flexibility.
Although we focus here on a particular form, welfare economics is quite flexible in the welfare functions it allows, the types of Pareto frontiers it admits, and the economic structures it supports (Adler, 2019). The formalism is also modular in its main building blocks, such as agents, resources, allocations, decisions, social planner, etc. This aligns well with the modular nature of learning objective in general (e.g., model architecture, loss function, regularization, learning algorithm). It also makes it easy to incorporate some aspects (and not others) when targeting a particular learning task.
Specification.
Accuracy as an objective is notoriously underspecified: for most learning tasks and model classes, there is typically a large set of (approximately) optimal models (Breiman, 2001; Marx et al., 2020; Hsu and Calmon, 2022). These all attain similarly high accuracy but can vary significantly in other important aspects, such as robustness (D’Amour et al., 2022), explainability (see Rudin, 2019), and fairness (e.g., Rodolfa et al., 2020; Coston et al., 2021; Black et al., 2022; Long et al., 2023). Welfare can be useful for breaking ties; in practice, this can sometimes be as simple as applying appropriate regularization (Levanon and Rosenfeld, 2021). We conjecture that in many applications it is possible to improve social outcomes with minimal drop in performance.
Interpretation.
Any learning objective encodes some implicit statement about social preferences; a welfare framework enables to uncover them. It also helps identify components missing in the objective (e.g., agency, interactions), highlight underspecified aspects, and critically examine possible outcomes.
Communication.
When speaking with practitioners, we have found it rather easy to convince that accounting for welfare is important, but quite challenging to explain how it can actually be done. Such efforts require an effective language for communicating novel concepts and complex ideas—which welfare economics provides. Communication is also the reason why our framework has supervised prediction at its core: it is fundamental to learning, widely familiar, and enables us to build around it increasingly complex economic constructs and at different levels of abstraction—namely our ‘orders’.
Impact.
Our final reason for focusing on welfare economics is its proven track record in connecting economic theory, practical econometrics, and real-world policymaking. Machine learning is renowned for its tight and fruitful relations between theory and practice—but has yet to establish an effective interface with policymakers and planners. A welfare perspective can help advance in this direction.
3.2 Alternative views (or: Why not welfare?)
The claim that machine learning should promote social welfare is unlikely to be contested; the debate lies in how to achieve this. Several subfields have made advancements towards particular aspects of this general objective, while others operate on the assumption that accuracy maximization suffices. We challenge these views, and assert that the field at large is still missing a holistic, unifying framework offering actionable guidelines. Here we highlight some alternatives and their connection to our position.
Give accuracy time.
One perspective is that if we give data enough time to accumulate and new methods enough time to improve, then predictive machine learning will organically overcome the concerns we have identified. One example to draw on is how despite many advances in optimization, the simple gradient descent algorithm still drives most modern tools. Another is how large language models have demonstrated that simply predicting the next word with sufficient accuracy and on enough data gives rise to emergent phenomena far beyond this basic task. Perhaps things will just work out?
We are not so optimistic. Our go-to example is the field of recommender systems: Despite a clear incentive to optimize welfare, recommendation systems have for decades settled on accurate prediction of item relevance as a proxy. At the time, and given the inherent difficulties that welfare presents, adopting the probability ranking principle (Robertson, 1977) as a guideline was a reasonable compromise. But over time, we are witnessing the negative implications of prediction-driven recommendations—on individual users, on suppliers and content creators, and on society. Surprisingly (or not), the same philosophy has now come to govern how LLMs are being fine-tuned; consider that methods like RLHF (Ziegler et al., 2019) and DPO (Rafailov et al., 2023) are similarly grounded on inferring user preferences. Predicting "what people want" addresses one important aspect of welfare, but falls short on others. As in recommendation, we conjecture that LLMs will give rise to various forms of scarcity, with competition or conflicts ensuing.
Divide and conquer.
Even if machine learning as a standalone solution does not suffice, one could argue that an economic approach can be applied on top of existing learning tools, rather than integrated within them as we propose. Hence, learning and policy can be advanced independently and combined only later. This is reasonable, and independent efforts and application will likely be required regardless of whether direct integration works or not. But there is increasing evidence that this will not suffice; in fact, the field of fair machine learning rose in response to the clear need for embedding social considerations within the learning objective itself. Advances in the study of fairness in learning have also shown that fairness constraints alone cannot guarantee equity, such as when the effects of learning accumulate over time (e.g., Liu et al., 2018) or as a result of user self-selection (Horowitz et al., 2024). We take these to suggest that interfacing learning and economics requires a holistic approach that enables to anticipate such issues, rather than addressing them separately and in hindsight.
Fairness.
Perhaps closest to our vision, the literature on fairness in machine learning (Dwork et al., 2012) explicitly aims to advance a certain form of social good, and for this it has received much attention. One concern that has been voiced is that fairness focuses on making things equal, but is unable to express the benefit that results from such efforts (Heidari et al., 2018; Hu and Chen, 2020; Kasy and Abebe, 2021). Some initial works have proposed welfare as an alternative (Shirali et al., 2024; Perdomo, 2024), but have yet to gain much influence. From our perspective, fairness takes one step towards acknowledging that something is of limited supply (typically positive predictions), but this is rarely made explicit. More importantly, fairness constraints do not account for agency: they do not ask what users want, nor how much they stand to lose or gain. Fairness constraints also do not anticipate how users will act; for example, predictions for school admissions may satisfy demographic parity, but if members of the minority group do not choose to apply initially, then this guarantee is broken. Other than some notable exceptions (e.g., Kannan et al., 2017), works on fairness generally do not consider (un)truthful reporting, manipulation, or information asymmetry—all central aspects of economic agent modeling. Whereas fairness typically aims to "solve" inequity, welfare provides a framework for advancing flexible goals.
Strategic learning.
Another relevant line of research is that of strategic learning, such as strategic classification (Brückner et al., 2012; Hardt et al., 2016) and performative prediction (Perdomo et al., 2020), which has drawn much recent interest. Strategic classification considers agency, but mostly in the form of “gaming”, and does not model interactions despite its canonical examples (e.g., admissions) having inherent constraints (e.g., quotas) (Liu et al., 2022). Some works discuss social aspects of learning (e.g., Milli et al., 2019), but the primary focus remains to be on maximizing accuracy. Performative prediction defines optimality and stability only w.r.t. the loss function, and so performatively optimal solutions can in principle have arbitrarily low welfare. Welfare is not a driving consideration, nor is it an integral part of current approaches. Focus in both fields has so far been predominantly theoretical, with little work done on practical tools and empirical analysis.
Synergy.
Welfare presents some inherent challenges for learning that may initially seem discouraging. We propose to draw inspiration from other fields that have successfully contended with similar or comparable challenges. Adversarial learning (Goodfellow et al., 2014) is a prime example for a field that has transformed an inherently game-theoretic problem to an engaging and highly relevant applied task. It’s worst-case perspective, however, can be very limiting, and the field may benefit from finer modeling of incentives and resources (Ehrenberg et al., 2025). Decision-focused learning (Mandi et al., 2024) and algorithms with predictions (Mitzenmacher and Vassilvitskii, 2022) have successfully built on predictive machinery to solve complex downstream tasks, as has the emerging field of prediction-powered inference (Angelopoulos et al., 2023) for improving statistical inference. A welfare framework can provide these fields with new economics-driven challenges. Both agency and resource limitations can often be cast as particular distribution shifts; distributional robustness (Chen et al., 2020a) and out-of-distribution generalization (Bühlmann, 2020) can provide tools, as well as benefit from new technical challenges.
Welfare without welfare.
Welfare economics is not the only approach for reasoning about and facilitating welfare, nor is it free from issues and limitations in itself. Criticism includes its subjective nature; the need to measure and compare utility across individuals; the emphasis on cardinal rather than ordinal utilities; the reliance on assumptions of rational behavior; the susceptibility to externalities and other sources of market failure; the need for a centralized social planner entity; and challenges in policy evaluation. Other schools of thought in economics offer alternatives: For example, Sen’s capabilities approach (Sen, 1999) focuses on ensuring people are capable of achieving what they seek, rather than the value of what they obtain. Within machine learning, there been have calls for alternative approaches as well, such as to ‘democratize’ the issue of alignment using social choice theory (Conitzer et al., 2024; Ge et al., 2024; Fish et al., 2024). We view these as complementary to ours, and believe there is merit in advancing welfare in machine learning simultaneously along several fronts.
3.3 From accuracy to welfare: paving the path
In courses and textbooks, machine learning is classically portrayed as a closed system: inputs go into a machine which then generates outputs. But even simple correlation machines that are deployed in the real world become part of it. For social settings comprising many other agents—namely us—we believe there is need for a refined systemic perspective. As the above examples convey, steps have been taken in various directions; these are praiseworthy, but lack unifying principles and guidance.
Although our goal is to discuss how learning can generally promote welfare, we believe there is value in keeping accuracy an integral part of the learning objective. This is both because accuracy has value in and of itself, and because predictive methods are, and will likely remain, a primary tool used in practice. We therefore argue for the deliberate extension of conventional supervised learning, not its replacement. Supervising learning already proposes various tools that can be harnessed, such as imposing cardinality constraints (Lapin et al., 2015; Petersen et al., 2022), controlling error types (Menon et al., 2019), or cost-sensitive optimization (Elkan, 2001). It also offers effective means for advancing general research such as benchmarks and competitions.
In Sec. 4 we present a proposed framework for welfare-aware learning that gradually extends from basic accuracy maximization towards increasingly complex economic objectives. One benefit of such a bottom-up framework is that it enables researchers and practitioners to choose the level of abstraction they wish to work in, such as assumptions on behavior, range of possible incentives, dependencies across agents, the role (or existence) of a central planner, and the type of pursued welfare objective. We believe this can make welfare-aware machine learning an engaging and appealing topic to work on for people of diverse backgrounds, interests, knowledge, and orientation.
Who drives welfare?
In the late 1940s, the field of economics found itself gaining significant social power, with its newly developed theoretical ideas and practical tools having unprecedented impact on society. Eighty years later, it appears we are witnessing similar trends for machine learning; it therefore only makes sense to try and learn from these earlier attempts—both successful and less so. One conclusion is that the wheel need not be (re)invented, only adapted. This is one of the grounds for our argument of drawing on ideas from welfare economics. But a major lesson learned is that economic objectives, welfare included, can rarely be forced upon a system. Economics views welfare as an emergent outcome arising from the decentralized interactions of many agents, particularly in idealized settings such as markets with perfect competition; this is the ‘invisible hand’ in essence. But in realistic setting, welfare needs guidance to materialize, whether through legislation, incentive alignment, or even minor nudges. As researchers, our goal should be to understand how systems can be steered towards beneficial outcomes, and what market failures can occur absent such direction. As system designers, we should carefully avoid coercing our own beliefs, and instead enable policymakers to implement their policies through the tools we build while maintaining transparency and accountability.
Who should care?
We believe the following different populations of researchers, with different levels of familiarity to welfare in machine learning, may all find our narratives helpful. The first, and perhaps foremost, is the general ML researchers and practitioners working on problems that have downstream social impact, [^3] Clear examples are people working on recommendation, social networks, and e-commerce. But we anticipate that over time, more and more subfields of ML will find themselves in a position where their work begins to have concrete social impact; consider text and image generation as recent examples. It is crucial to think one step ahead before we start to face serious welfare issues. but who do not account for this explicitly. Here our goal is not only to raise awareness to welfare issues, but also to provide a tangible means towards achieving it. This is why our framework below has supervised prediction at its core: it is fundamental to ML, widely familiar, and enables us to build around it increasingly complex economic constructs and at different levels of abstraction–namely our ‘orders’. The second population features researchers in related sub-fields such as fairness (Dwork et al., 2012) or strategic classification (Hardt et al., 2016), which is a considerably smaller group within the entire ML community. Here we intend to provide a broader framework as well as complementary perspective to the questions they are already pursuing. For example, fairness research has mostly overlooked agency, whereas strategic classification rarely considers resources. By revealing such “blind spots” we wish to motivate researchers to confront them. Finally, the third population consists of researchers already working at the intersection of ML and (welfare) economics. This is perhaps an even smaller group within the general ML community (for which utility and welfare have been overlooked). Our goal here is to provide an organizing conceptual framework that is firmly grounded in the core concepts of machine learning.
4 Proposed framework
@@
To concretize our position, we propose a formal framework for welfare-maximizing learning, organized into three hierarchical ‘orders’. Our framework presents a hierarchy of problem formulations for welfare-maximizing learning, organized into three ‘orders’ of gradually increasing complexity. The orders are built in a bottom-up fashion: The lowest order (Order 0) coincides with the standard objective of maximizing predictive accuracy, but provides a novel welfare perspective suited for social tasks. Each higher order then builds on and generalizes the one below it, this by adding to the objective another layer of economic complexity: first focusing on system decisions (Order 1), and then on user choices (Order 2). As such, our framework is based on accuracy maximization at its core, but enables—and in fact requires—to explicitly model resources, allocations, and agency. It also requires to specify the role predictions play in shaping social outcomes. An illustration of the framework and its orders is given in Fig. 1.
We view our framework as a means to enable the effective communication of ideas regarding welfare to different target audiences. For example, we have found it rather easy to convince ML practitioners that welfare is important, but quite challenging to explain how this can be done in practice. We envision our framework as enabling the flow of ideas and methodologies between ML and welfare economics in both directions: Economics offers many compelling ideas, but interfacing them with ML is rarely straightforward. Conversely, ML offers many powerful tools, but these have yet to be harnessed for maximizing welfare.
@@ 4.1 Order 0: Allocating Accuracy
Consider first the standard accuracy objective in Eq. (2), and its interpretation as a welfare objective in which users gain utility from obtaining accurate predictions. Accuracy will be beneficial to users in any a platform or service that relies on predictions; consider recommendation systems, online market platforms, financial aid tools, or diagnistic health services. We argue that in such cases, and as long as learning is prone to some level of error (which is plausible), then accuracy itself is a limited resource—simply because not all users can obtain the same level of accuracy. Thus, the choice of classifier becomes, in effect, a decision on which users will be ‘allocated’ more accurate predictions, and which less. This then begs the question: who should be allocated an accurate prediction? One answer is users who most need it, or who benefit mostly. In the reasonable case where there is natural variation in individual utilities, we can to optimize a utility-weighted accuracy objective:
@@
Maximizing welfare through learning presents new algorithmic challenges. One task is to identify learners that optimize a given social welfare function—intrinsically, a question of identifying a desirable point on the Pareto front. Another task is to generate all points on the frontier (see, e.g., Navon et al. (2021)). One possible approach is to cast Eq. (3) as a problem of distribution shift. Here shift can stem from several factors, including the welfare function weights, user utilities, and user inputs. But agency means that utilities can be misreported and features manipulated; as a result, the way in which the distribution shifts becomes dependent on the choice of classifier, indirectly through how users respond. This makes Eq. (3) a challenging instance of model-dependent distribution shift (Drusvyatskiy and Xiao, 2023). Here the interaction between the system and its users can be modeled as a Stackelberg game (Chen et al., 2020b). This perspective has been useful for questions on learnability, e.g. via generalization bounds that rely on strategic VC analysis (Zhang and Conitzer, 2021; Sundaram et al., 2023) or that accommodate feature-dependent utilities (Pardeshi et al., 2024). Nonetheless, many important open questions remain.
@@ 4.2 Order 1: Incorporating System Decisions
Generally, predictions are useful for the system if they aid in making better decisions about uncertain outcomes. Since decisions in social settings often directly prescribe how to allocate limited resources to users, a useful next step is to encode them explicitly into the objective. We formalize this idea by modeling a system that makes decisions about users (e.g., who to hire) through predictions (e.g., resume screening). Consider a user $x$ with label $y$, and denote by $a$ the action the system takes (e.g., hire or not). Let ${r}(a,y)$ be the reward for the system on this user given action $a$ (e.g., the quality of the candidate, if hired). The direct dependence of ${r}$ on $y$ means that if we know $y$, then we can write the optimal action as $a^{*}={\pi}(y)$ for some policy ${\pi}$. Since $y$ is generally unknown, the common approach is to replace $y$ in ${\pi}$ with a prediction ${\hat{y}}=h(x)$, denoted $a={\pi}(h(x))={\pi}_{h}(x)$, in hopes that better predictions translate to better decisions. Such policies, referred to as prediction policies (Kleinberg et al., 2015), are appropriate when the uncertainty in $y$ is a stronger factor for outcomes than potential causal effects (we discuss this distinction further in Section 5.4). In this setting, choosing to work with prediction policies gives reason for the system to optimize $h$ for accuracy (Eq. (2)).